Twice the larger of two numbers is three more than five times the smaller, and the
sum of four times the larger and three times the smaller is 72. What are the
numbers ?

Found 2 solutions by jorel555, MathTherapy:
Answer by jorel555(1290)

Let m and n be your two numbers.
4m+3n=72
4m=72-3n
m=18-3n/4
2(18-3n/4)-3=5n
8(18-3n)-12=20n
144-24n-12=20n
132=44n
n=3
m=63/4!!!!!!!!!!!!

Answer by MathTherapy(10556)


Twice the larger of two numbers is three more than five times the smaller, and the
sum of four times the larger and three times the smaller is 72. What are the
numbers ?